1)  What is the solubility of PbI2 in g/l, if the solubility product is 7.9x10-9

2) Cu(s)/ Cu+2 (0.02M) //Zn2+ (0.02M)/Zn(s); the standard reduction potential for Cu and Zn is 0.34 and -0.76 respectively. Calculate the standard cell potential, change in free energy and cell potential


1
Expert's answer
2021-05-08T23:33:36-0400

1.) Let the solubility of "PbI_2" in 0.01M NaI solution be s M.

Then in this case,

"[Pb^{2+}] = sM"

And,

"[I^{-}] = (2s+0.01)M"

So, "K_{sp} = [Pb^{2+}][I^{-}]^2"

"1.4 \times 10^{-8} = s \times (4s^2+4s(0.01)+(0.01)^2)"

Neglecting "s^2" and "s^3" terms, we get:

We get,

"s = 1.4 \times 10^{-4}M"

2.) In this part we can use the Nernst Equation,


"E_{cell} = E_0 - \dfrac{0.0591}{n}log\dfrac{[Reduced]}{[Oxidised]}"